Mrs. Alan travels from home to work at 60 km per Hour (Km/h).
She returns along the same route. Her Average speed for the round trip is 72 km/h, What is her speed from work to homes, in km/h?

Found 2 solutions by mananth, songtao:
Answer by mananth(16949)
Answer by songtao(1)

(60+x)/2=72, 60+x=72x2=144, x=144-60=84. So, her speed from work to homes in 84km/h.